using the equilibrium liners to find a Qd and Qs equation.
S. 7 2 -7Q 16 20 De Q 20-2P
Answer #1

Solution

From the given figure above;

The demand and supply curves interesect at price level equal to 7.

Therefore, equilibrium price is equal to 7

At the equilibrium,Demand =Supply

Q = 20-2p = 20-2*7=20-14=6

Therefore; Q=S=6 at the equilibrium